# ERX: A Fast Real-Time Anomaly Detection Algorithm for Hyperspectral Line-Scanning

### v1

@@ II-E Dimensionality Reduction
where $\bm{X}_{t}$ is the most recent hyperspectral line and $\bm{S}_{t}$ is the line after the dimensionality reduction, with the eigenvalues ${e}_{t}$ forming the diagonal of the lower-dimensional covariance matrix. The Mahalanobis distance was calculated as:
$\delta_{i}$ $=\sqrt{\bm{s}_{i}\text{diag}(\bm{e}_{t})^{-1}\bm{s}_{i}^{T}}$ (13)
where $\bm{s}_{i}$ is the $i$th dimensionality-reduced pixel in the current line. The authors introduced an adaptive threshold method to detect anomalous pixels by excluding their contribution to background statistics.
LbL-FAD was also proposed, using a modified Gram-Schmidt method for orthogonal subspace projection [33]. This algorithm has also been adapted for hardware through field-programmable gate arrays (FPGAs), demonstrating the efficiency of the algorithm onboard small platforms in real-time [34, 35].
@@ III Proposed Algorithm - ERX
The Exponentially moving RX detector (ERX) uses exponentially moving averages to perform rapid updates to the background mean and covariance while keeping memory requirements low. The use of exponential weightings allows for a combination of global and local anomaly detection.
@@
For a new hyperspectral line collected $\bm{X}_{t}\in\mathbb{R}^{p,b}$, the mean vector and covariance matrix are estimated:
$\bm{\bar{x}}_{t}$ $=\frac{1}{p}\sum\limits_{i=1}^{p}\bm{x}_{i}$ (14)
$\bm{\bar{K}}_{t}$ $=\frac{1}{p-1}\sum\limits_{i=1}^{p}(\bm{x}_{i}-\bm{\bar{x}}_{t})(\bm{x}_{i}-\bm{\bar{x}}_{t})^{T}$ (15)
@@
where each pixel in $\bm{X}_{t}$ is defined as $\bm{x}_{i}\in\mathbb{R}^{b}$ ($b$ is the number of bands, $p$ is the number of pixels per line, and $i=1,2,...,p$). The background mean and covariance are then updated using a momentum factor $\alpha$ (where $10<\alpha<$ 1):
$\bm{\mu}_{t}$ $=(1-\alpha)\bm{\mu}_{t-1}+\alpha\bm{\bar{x}}_{t}$ (16)
$\bm{K}_{t}$ $=(1-\alpha)\bm{K}_{t-1}+\alpha\bm{\bar{K}}_{t}$ (17)
Increasing $\alpha$ weighs recent lines more heavily, and decreasing $\alpha$ puts more weight on the initial background statistics. With the background statistics calculated, the Mahalanobis distance each pixel is defined as:
$\delta_{i}$ $=\sqrt{(\bm{x^{\prime}}_{i}-\bm{\mu}_{t})\bm{K}_{t}^{-1}(\bm{x^{\prime}}_{i}-\bm{\mu}_{t})^{T}}$ (18)
where $\bm{x^{\prime}}_{i}$ is each pixel in the previously captured line $\bm{X}_{t-o}$, $o$ being a small offset. Calculating the Mahalanobis distance of a previous line allows ERX to incorporate spatial information ahead of the pixel when estimating the background mean and covariance. Like Zhang et al. [22], Cholesky decomposition is used on the covariance matrix, giving:
$\delta_{i}$ $=\sqrt{[(\bm{x^{\prime}}_{i}-\bm{\mu}_{t})\bm{L}_{t}^{-1}][(\bm{x^{\prime}}_{i}-\bm{\mu}_{t})\bm{L}_{t}^{-1}]^{T}}$ (19)
where $\bm{L}_{t}$ is the lower triangle of $\bm{K}_{t}^{-1}+10^{-5}\bm{I}$. The small identity matrix $10^{-5}\bm{I}$ is added to the covariance matrix prior to Cholesky decomposition to prevent matrix singularities and improve the stability of the algorithm. (19) can be simplified by substituting $\bm{z}_{i}=(\bm{x^{\prime}}_{i}-\bm{\mu}_{t})\bm{L}_{t}^{-1}$:
$\delta_{i}$ $=\sqrt{\bm{z}_{i}\cdot\bm{z}_{i}^{T}}$ (20)
where $\bm{z}_{i}$ is calculated directly using forward substitution for $\bm{L}_{t}\bm{z}_{i}=(\bm{x^{\prime}}_{i}-\bm{\mu}_{t})$, avoiding the inversion of the covariance matrix. The Mahalanobis distance vector is then normalised across the line:
@@
where $\bar{\delta}$ is the mean Mahalanbois distance and $\sigma_{\delta}$ is the standard deviation. Pixels in the line are defined as an anomaly if they are above a threshold value $T_{N}$:
$\hat{y}_{i}$ =$=\begin{cases}1,&\text{if }\delta_{N,i}\geq T_{N}\\ 0,&\text{otherwise}\end{cases}$ (22)
An optimised pseudo-code version of the algorithm is detailed in algorithm 1, with vector notation and broadcasting assumed to speed up processing.
Algorithm 1 ERX Algorithm
Input: $\bm{X}_{t}$: current line matrix, $\bm{\mu}_{t-1}$: previous mean vector, $\bm{K}_{t-1}$: previous covariance matrix, buffer: rolling buffer with $l$ previous lines, $\alpha$: momentum value, $o$: line offset, $n_{l}$: the total number of lines processed so far, $T_{N}$: normalised threshold value.
Output: $\bm{y}_{t}$: binary vector of anomaly detections in line, $\bm{\mu}_{t}$: new mean vector, $\bm{K}_{t}$: new covariance matrix, buffer: updated buffer.
1:procedure ProcessLine($\bm{X}_{t}$, $\bm{\mu}_{t-1}$, $\bm{K}_{t-1}$, buffer, $\alpha$, $o$, $n_{l}$, $T_{N}$)
2:     $\bm{\bar{x}}_{t}\leftarrow\text{mean}(\bm{X}_{t})$
3:     $\bm{\bar{K}}_{t}\leftarrow\text{covariance}(\bm{X}_{t})$
4:     $\textbf{buffer}\leftarrow\textbf{buffer}\text{.append}(\bm{X}_{t})$
5:     $\textbf{buffer}\leftarrow\textbf{buffer}\text{.remove}(\bm{X}_{t-l})$
6:     if $n_{l}=0$ then
7:         $\bm{\mu}_{t}\leftarrow\bm{\bar{x}}_{t}$
8:         $\bm{K}_{t}\leftarrow\bm{\bar{K}}_{t}$
9:     else
10:         $\bm{\mu}_{t}\leftarrow\bm{\mu}_{t-1}-\alpha(\bm{\mu}_{t-1}-\bm{\bar{x}}_{t})$
11:         $\bm{K}_{t}\leftarrow\bm{K}_{t-1}-\alpha(\bm{K}_{t-1}-\bm{\bar{K}}_{t})$
12:     if $n_{l}>=l$ then
13:         $\bm{X}_{t-o}\leftarrow\textbf{buffer}[t-o]$
14:         $\bm{L}_{t}\leftarrow\text{cholesky}(\bm{K}_{t}+10^{-5}\bm{I}_{b})$
15:         $\bm{\Delta}_{t}\leftarrow\bm{X}_{t-o}-\bm{\mu}_{t}$
16:         $\bm{z}_{t}\leftarrow\text{solve}(\bm{L}_{t},\bm{\Delta}_{t}^{T})$
17:         $\bm{\delta}_{t}\leftarrow\text{norm}(\bm{z}_{t})$
18:         $\bm{\delta}_{N}\leftarrow\frac{\bm{\delta}_{t}-\text{mean}(\bm{\delta}_{t})}{\text{variance}(\bm{\delta}_{t})}$
19:         $\bm{\hat{y}}_{t}\leftarrow\text{array}(\text{len}(\bm{\delta}_{N}))$
20:         for $i=1$ to len($\bm{\hat{y}}_{t}$do
21:              if $\bm{\delta}_{N}[i]>=T_{N}$ then
22:                  $\bm{\hat{y}}_{t}[i]\leftarrow 1$
23:              else
24:                  $\bm{\hat{y}}_{t}[i]\leftarrow 0$
25:     $n_{l}\leftarrow n_{l}+1$
26:     return $\bm{y}_{t}$, $\bm{\mu}_{t}$, $\bm{K}_{t}$, buffer, $n_{l}$
IV Experiments & Results
This section includes the datasets, experimental designs, and the results. The Python code for all algorithms and their experiments is available at https://github.com/WiseGamgee/HyperAD.
@@ IV-A Datasets
The beach dataset, sized 452 pixels by 3072 lines with 108 bands, was collected via a drone-mounted hyperspectral line-scan system on a beach in Queensland, Australia. It transitions from vegetation, to sand, to shallow ocean, and includes anomalous objects such as different coloured and sized tarps, two vehicles, and reflective cones (the smaller yellow objects in the ground truth). This version uses radiance and lacks geometric correction, as is evident from the ”wobble” between lines in Figure 4. For more details, see Mao et al. [36].
@@
The synthetic dataset, sized 600 pixels by 2400 lines with 90 bands, was created by sampling pixels from an AVIRIS radiance dataset collected near Gulfport, Mississippi [37]. The background is generated from vegetation, sand, and water pixels with smoothed transitions between each type. The anomalies are square targets made from airport runway pixels. The targets are equally spaced with decreasing size along-track (left to right), and are repeated from top to bottom with 0%, 10%, 20%, 30%, and 50% background mixing to add noise. The original 224 bands are reduced to 189 by removing the water absorption and low signal-to-noise ratio bands (as in Głomb and Romaszewski [38]). After this, only the first 90 bands are kept to further improve the signal-to-noise ratio and avoid the OpenBLAS threading issue (discussed in Section V-A).
@@
The Sequoia National Park (SNP) dataset, 1116 pixels by 2499 lines with 13 bands, was collected using the Sentinel 2 multispectral satellites. The dataset is of a section of the Sequoia National Park in California, U.S.A and includes live wildfires as anomalies which were identified using the infrared bands. It uses level-1c processing for top-of-atmosphere reflectance, unlike the hyperspectral datasets that use radiance. This multispectral dataset helps demonstrate algorithm performance on lower dimensional data.
@@ IV-B Anomaly Detection Algorithms
The following algorithms are tested on the datasets:
@@
  5. 5.
    ERX, presented in algorithm 1.
TABLE I: Normalised Mahalanobis Distance Threshold Values ($T_{N}$) for Every Algorithm and Dataset.
Algorithm Beach Dataset Synthetic Dataset SNP Dataset
RX Baseline 1.5 3 5
RT-CK-RXD 1.5 3 5
RX-BIL 1.5 3 5
LBL-AD 1.5 1.5 4
ERX 1.5 3 5
@@
The RX Baseline algorithm serves as a benchmark for real-time line-scan anomaly detection. A rolling buffer stores the last 99 lines. The mean vector and covariance matrix are calculated using the whole buffer, and the RX algorithm is applied to the centre line. This process is repeated for each new line (visualised in Figure 7).
Figure 7: Visualisation of the RX Baseline algorithm. This process is repeated for each new line captured by a line-scan camera in real-time.
RT-CK-RXD uses a buffer of length 99 for the initial mean and covariance estimates. This algorithm updates the mean and covariance and computes the Mahalanobis distance pixel-wise. Just-in-time compilation via Numba [39] is used to speed up inverse covariance updates.
@@
RX-BIL also uses a 99 line buffer and randomly discards 50% of the pixels per line. Instead of initialising the mean vector and correlation matrices with small random numbers, the mean and correlation of the first line are used for initialisation, substantially speeding up convergence.
For LBL-AD, the bands are reduced to 3 principal components. Horstrand et al. [14] recommend a maximum of five principal components to avoid impacting speed too much. The automatic threshold component is not included, and Numba is used to speed up the power iteration algorithm.
ERX uses a buffer of length 99, an offset of $o=30$, and a momentum factor of $\alpha=0.5$. In practice, each algorithm needs a threshold selected to detect anomalous pixels. Optimal normalised thresholds (22) for each algorithm and dataset are determined by grid search and shown in Table I. The preset thresholds are used to demonstrate potential anomaly detection performance, but are noted as limiting due to needing prior information.
IV-C Experiments
The first experiment evaluates the impact of varying the ERX momentum parameter $\alpha$. Two further experiments compare ERX’s speed and detection performance to the other algorithms. All experiments were carried out on an NVIDIA Jetson Xavier NX compute module with 8GB RAM, a 6-core NVIDIA Carmel ARMv8.2 CPU, and a 384-core NVIDIA Volta GPU with 48 Tensor cores, using the default 10W power mode.
IV-C1 ERX Momentum Evaluation
This experiment investigates the ERX momentum parameter $\alpha$, which modifies the algorithm’s detection performance based on anomaly size. The Area Under the ROC Curve (AUC) is used to assess performance across a range of thresholds $T_{N}$, and the F1 score is used because it is more robust than the AUC for highly imbalanced datasets. The thresholds used to calculate the F1 scores are defined in Table I. The results for each dataset are given in Table II. Additionally, binary anomaly maps of the synthetic dataset in Figure 8 demonstrate how momentum affects ERX’s detection performance.
TABLE II: ERX Detection Results - Varying Momentum ($\alpha$).
Beach Dataset Synthetic Dataset SNP Dataset
$\alpha$ F1 AUC F1 AUC F1 AUC
0.9 0.463 0.848 0.157 0.855 0.313 0.986
0.5 0.441 0.845 0.150 0.852 0.306 0.986
0.1 0.310 0.809 0.085 0.633 0.279 0.987
0.01 0.302 0.793 0.085 0.547 0.303 0.989
1e-3 0.369 0.801 0.060 0.503 0.397 0.992
1e-4 0.439 0.796 0.063 0.523 0.483 0.993
1e-5 0.450 0.789 0.070 0.530 0.444 0.993
((a))
((b))
((c))
((d))
Figure 8: ERX’s binary anomaly detection maps for various momentum values when applied to the synthetic dataset.
IV-C2 Speed Comparison

### v2

@@ II-F Dimensionality Reduction
where $\bm{X}_{t}$ is the most recent hyperspectral line and $\bm{Z}_{t}$ is the line after the dimensionality reduction, with the eigenvalues ${e}_{t}$ forming the diagonal of the lower-dimensional covariance matrix. The Mahalanobis distance was calculated as:
$\delta_{i}$ $=\sqrt{\bm{z}_{i}\text{diag}(\bm{e}_{t})^{-1}\bm{z}_{i}^{T}}$ (13)
where $\bm{z}_{i}$ is the $i$th dimensionality-reduced pixel in the current line. The authors introduced an adaptive approach that prevents anomalous pixels from contributing to the background statistics. The authors recommend reducing the bands to a maximum of five principal components, as any more significantly slows down processing.
LbL-FAD was also proposed, using a modified Gram-Schmidt method for orthogonal subspace projection [64]. This algorithm has also been adapted for hardware through field-programmable gate arrays (FPGAs), demonstrating the efficiency of the algorithm onboard small platforms in real time [69, 70]. These dimensionality reduction algorithms are fast, accurate, and scalable. However, like the vast majority of related work, they are typically evaluated on small, uniform, and geometrically corrected HSI datasets. These datasets do not reflect the changing scenery and distortions encountered by drones or satellites when surveying large areas in real time.
II-G Summary
Existing algorithms used a series of methods to enhance the RX algorithm and adapt it for real-time line scanning. These were grouped into However, these algorithms lacked the combination of four key characteristics: fast speeds for onboard implementation, scalability to a large number of HSI bands, adaptability to changing scenery, and robustness to geospatial and radiometric distortions in real-time data. In addition, the previous work used only on small, geometrically corrected image datasets, with little variation in scenery. Thus, there is a gap in assessing the true adaptability and robustness of line-scanning algorithms under more practical conditions.
In the following section, we present ERX and three larger and more diverse datasets to address these challenges. Two datasets have changing scenery to test adaptability, with one also having geometric distortions to better evaluate robustness.
@@ III Proposed Algorithm - ERX
III-A Overview
Figure 4: ERX algorithm flowchart. SRP is sparse random projection, and $\bm{\hat{\mu}}_{t}$ and $\bm{\hat{K}}_{t}$ are the mean and covariance of the current line. $\bm{\mu}_{t}$ and $\bm{K}_{t}$ are the background mean and covariance, and EMA refers to the exponentially moving averages. $\bm{\delta}_{t}$ is the Mahalanobis distance vector of all pixels in the line.
The Exponentially moving RX algorithm (ERX) is an online algorithm, which means that it detects anomalies while simultaneously updating itself in real time. Figure 4 illustrates the sequential processing of hyperspectral lines by ERX, one line at a time. ERX features two main components: sparse random projection (SRP) and exponentially moving averages (EMAs). SRP enhances scalability by reducing the number of hyperspectral bands to a smaller number of dimensions. This significantly enhances the processing speed, as the number of bands is a major performance bottleneck for real-time deployment. Line-by-line EMAs rapidly update the background mean and covariance, prioritising recent lines and adapting to changing scenery.
Initially, a small buffer period is used to improve the modelling of the background mean and covariance before starting detection. For this research, the buffer is set to 99 lines, but in practice ERX can start detection immediately without storing previous hyperspectral lines. Cholesky decomposition and forward substitution are used to calculate the Mahalanobis distance of each pixel, with minor stability improvements over conventional methods. The algorithm concludes with a line-by-line normalisation of the Mahalanobis distances to identify anomalous pixels. This process is repeated for any number of hyperspectral lines. ERX’s Python implementation is openly available under “detectors” in the GitHub repository. The remainder of this section expands on the main components of ERX.
III-B Sparse Random Projection (SRP)
The large number of hyperspectral bands is one of the biggest drags on performance and scalability for real-time anomaly detection. Increasing the number of spectral bands disproportionately decreases the processing speed and exposes algorithms to the Hughes phenomenon [71], also known as the curse of dimensionality. This means that the pixels are more spread out in the feature space, making the separation of anomalies from the background more difficult. To alleviate these issues, ERX uses sparse random projection (SRP) [72, 73] to reduce the original hyperspectral bands to a smaller number of dimensions.
Random projection reduces the input dimensionality using a matrix of randomly generated weights [74]. SRP zeroes many of these weights to further reduce the computational load. SRP is simpler and more efficient compared to other methods such as PCA [75], and still preserves the relative distances between the original data. SRP has been used to reduce the dimensionality of text and image data [74], but this work appears to be the first use in hyperspectral anomaly detection.
In the context of ERX, the random weight matrix $\bm{W}\in\mathbb{R}^{b,d}$ is generated during initialisation with the following distribution ($b$ is the number of bands and $d$ is the new number of dimensions after projection):
$w_{i,j}$ $=\frac{\sqrt{s}}{\sqrt{d}}\begin{cases}1,&\text{with probability }\frac{1}{2s}\\ 0,&\text{with probability }1-\frac{1}{s}\\ -1,&\text{with probability }\frac{1}{2s}\end{cases}$ (14)
where $w_{i,j}$ is any weight in $\bm{W}$, and the sparsity parameter $s=\sqrt{b}$ [73]. The number of active (non-zero) weights in the random projection matrix equals the number of matrix elements multiplied by the activation probability:
$n_{aw}$ $=(b\times d)\times\frac{1}{\sqrt{b}}$ (15)
$=d\sqrt{b}$ (16)
Given the latest hyperspectral line $\bm{X}_{t}\in\mathbb{R}^{p,b}$ (where $b$ is the number of bands, $p$ is the number of pixels per line), every pixel is projected to the lower number of dimensions ($d$):
$\bm{Z}_{t}=\bm{X}_{t}\cdot\bm{W}$ (17)
where $\bm{Z}_{t}\in\mathbb{R}^{p,d}$ is the reduced hyperspectral line with $d$ dimensions.
The Johnson-Lindenstrauss Lemma is typically used to determine the smallest number of dimensions [72] that preserves the original distances between the spectral signatures of the original HSI. Unfortunately, this approach requires an initial image sample to estimate the optimal $d$, which is not available for this task. Hence, $d$ is empirically set as $5$ for all tests, with the effect of altering $d$ explored in Section V-C.
Empirical dimension selection also improves flexibility by removing the dependency on an initial background sample, enabling adaptation to changing backgrounds. This contrasts with the existing algorithms that use dimensionality reduction methods, which assume a uniform or constant background. This implementation of SRP improves processing speed, reduces memory usage, and boosts the signal-to-noise ratio without any optimisation. However, random feature selection introduces some variability in model performance, which is investigated through repeated experiments. The SRP implementation of the scikit-learn library is used [76].
III-C Exponentially Moving Mean & Covariance
Existing algorithms, such as those that use the Woodbury matrix identity, update the background statistics incrementally or through methods such as simple moving averages (SMAs). Incremental updates can be fast, but treat all pixels equally which limits adaptability to scene changes. SMAs provide adaptability by focusing on a smaller window of pixels, though they operate more slowly. ERX uses exponentially moving averages (EMAs) to rapidly update the background mean and covariance estimates, line-by-line. This fast operation prioritises recent lines, directly enhancing adaptability.
This concept is demonstrated in Figure 5, where the use of an exponential momentum factor ($\alpha$) is used to change the “memory” of the algorithm. Increasing the momentum parameter $\alpha$ increases the weighting of recent lines, shortening ERX’s memory and making older lines less relevant for anomaly detection. This allows ERX to adapt to changing scenery in real time. This seems to be the first use of line-by-line exponentially moving statistics for HAD.
Figure 5: Conceptual comparison of ERX’s exponentially moving averages (EMAs), versus equal weighting. EMAs prioritise newer lines and forget older ones (top), unlike equally weighted methods, which treat all lines uniformly (middle). The plot (bottom) displays the weight profiles for all lines in the example scene, with the most recent captured at time $t$.
@@ III-C Exponentially Moving Mean & Covariance
To implement this, the mean vector ($\bm{\hat{\mu}}_{t}\in\mathbb{R}^{d}$) and covariance matrix ($\bm{\hat{K}}_{t}\in\mathbb{R}^{d,d}$) of the projected hyperspectral line ($\bm{Z}_{t}$) are calculated:
$\bm{\hat{\mu}}_{t}$ $=\frac{1}{p}\sum\limits_{i=1}^{p}\bm{z}_{i}$ (18)
$\bm{\hat{K}}_{t}$ $=\frac{1}{p-1}\sum\limits_{i=1}^{p}(\bm{z}_{i}-\bm{\hat{\mu}}_{t})(\bm{z}_{i}-\bm{\hat{\mu}}_{t})^{T}$ (19)
@@
where each pixel in $\bm{Z}_{t}$ is defined as $\bm{z}_{i}\in\mathbb{R}^{d}$ ($d$ is the number of dimensions after SRP, $p$ is the number of pixels per line, and $i=1,2,...,p$). The background mean ($\bm{\mu}_{t}$) and covariance ($\bm{K}_{t}$) are then updated using the momentum factor $\alpha$ (where $10<\alpha<$ 1):
$\bm{\mu}_{t}$ $=(1-\alpha)\bm{\mu}_{t-1}+\alpha\bm{\hat{\mu}}_{t}$ (20)
$\bm{K}_{t}$ $=(1-\alpha)\bm{K}_{t-1}+\alpha\bm{\hat{K}}_{t}$ (21)
The default momentum value is empirically set to $\alpha=0.1$ for most experiments, and the initial background values when $t=0$ are given by the mean and covariance of the first line. Momentum can affect ERX’s detection performance and is examined in detail in Section V-C.
III-D Cholesky Decomposition & Improved Stability
Like Zhang et al. [66], Cholesky decomposition is used to quickly compute the Mahalanobis distance by avoiding matrix inversion. Due to the correlation matrix’s instability and singularity risk, ERX simply uses the covariance matrix $\bm{K}_{t}^{-1}$ plus a small identity matrix $10^{-5}\bm{I}$ instead:
$\delta_{t}$ $=\sqrt{[(\bm{z}_{i}-\bm{\mu}_{t})\bm{L}_{t}^{-1}][(\bm{z}_{i}-\bm{\mu}_{t})\bm{L}_{t}^{-1}]^{T}}$ (22)
where $\bm{L}_{t}$ is the lower triangle of $\bm{K}_{t}^{-1}+10^{-5}\bm{I}$. Adding a small identity matrix enhances the stability of the algorithm by ensuring a positive definite covariance matrix, resolving failures found in previous algorithms. (22) can be simplified by substituting $\bm{m}_{i}=(\bm{z}_{i}-\bm{\mu}_{t})\bm{L}_{t}^{-1}$:
$\delta_{i}$ $=\sqrt{\bm{m}_{i}\cdot\bm{m}_{i}^{T}}$ (23)
where $\bm{m}_{i}$ is calculated directly using forward substitution for $\bm{L}_{t}\bm{m}_{i}=(\bm{z}_{i}-\bm{\mu}_{t})$. This approach provides a more robust solution when using Cholesky decomposition.
III-E Line-by-line Normalisation
Using a threshold is the final step in anomaly detection, distinguishing Mahalanobis distances above the boundary as anomalies and those below as normal. ERX first normalises the Mahalanobis distance vector for better anomaly-background separation:
@@ III-E Line-by-line Normalisation
where $\bar{\delta}$ is the mean Mahalanobis distance and $\sigma_{\delta}$ is the standard deviation. Pixels are labelled anomalies if they exceed the threshold $\tau_{N}$:
$\hat{y}_{i}$ =$=\begin{cases}1,&\text{if }\delta_{N,i}\geq\tau_{N}\\ 0,&\text{otherwise}\end{cases}$ (25)
IV Experimental Setup
This section includes the datasets, experimental design, and the results. The datasets and the Python code for all algorithms and their experiments are available at https://github.com/WiseGamgee/HyperAD.
@@ IV-A Datasets
The beach dataset, sized 452 pixels by 3,072 lines with 108 bands, was collected via a drone-mounted hyperspectral line-scan system on a beach in Queensland, Australia. The resolution of the pixels is 5 cm. It transitions from vegetation, to sand, to shallow ocean, and includes anomalous objects such as different coloured and sized tarps, two vehicles, and reflective cones (the smaller yellow objects in the ground truth). This version uses radiance and lacks geometric correction, as is evident from the “wobble” between lines in Figure 6. For detailed dataset collection methods, see Mao et al. [15].
@@
The synthetic dataset, sized 600 pixels by 2,400 lines with 90 bands, was created by sampling 7.1m pixels from an AVIRIS radiance dataset collected near Gulfport, Mississippi [77]. The background is generated from vegetation, sand, and water pixels with smoothed transitions between each type. The anomalies are square targets made from airport runway pixels. The targets are equally spaced with decreasing size along-track (left to right), and are repeated from top to bottom with 10%, 20%, 30%, and 50% background mixing to add noise. The original 224 bands are reduced to 189 by removing the water absorption and low signal-to-noise ratio bands (as in Głomb and Romaszewski [78]). After this, only the first 90 bands are kept to further improve the signal-to-noise ratio and avoid the OpenBLAS threading issue (discussed in Section V-A).
@@
The Sequoia National Park (SNP) dataset, 1,116 pixels by 2,499 lines with 13 bands, was collected using the Sentinel-2 multispectral satellites and downloaded via Sentinel Hub. It has a 30m pixel resolution. The dataset covers part of Sequoia National Park in California and includes live wildfires as anomalies that were identified using infrared bands. It uses L1C processing for top-of-atmosphere reflectance, unlike the hyperspectral datasets that use radiance. This multispectral dataset helps to demonstrate algorithm performance on less correlated lower-dimensional data.
@@ IV-B Anomaly Detection Algorithms
The following algorithms are compared in the experiments:
@@
ERX uses a 99 line buffer to estimate the exponentially moving mean and covariance before detection starts. A momentum factor of $\alpha=0.1$ is used, and the bands are reduced to 5 dimensions via sparse random projection (that is, $d=5$) for all experiments. The RX Baseline algorithm serves as a benchmark for real-time anomaly detection. A rolling buffer stores the most recent 99 lines. The mean vector and covariance matrix are calculated using the entire buffer, and the RX algorithm is applied to the centre line. This process is repeated for each new line.
@@
RT-CK-RXD detects anomalies using a pixel-wise Woodbury matrix identity implementation. A buffer of 99 lines is used for the initial mean and covariance estimates. Just-in-time compilation via Numba [79] is also used to speed up inverse covariance updates. RX-BIL uses a line-by-line Woodbury matrix identity implementation, as well as the correlation matrix instead of the covariance matric. A 99 line buffer is also used and $\eta=50\%$ of the pixels per line are randomly discarded. Instead of using small random numbers to initialise the mean vector and correlation matrices as described in the paper, the mean and correlation of the first line are used for initialisation, substantially speeding up convergence.
LBL-AD enhances scalability using real-time PCA, offering a good comparison for ERX’s dimensionality reduction via SRP. The bands are reduced to three principal components ($d=3$). Numba is used to speed up the power iteration algorithm. The implementations of all algorithms are available on the GitHub repository.
IV-C Metrics
The speed of each algorithm is simply measured as the average number of lines processed per second (LPS) for a given dataset. The detection performance of each algorithm is measured primarily by using the average receiver operating characteristic (ROC) curve. The ROC curve plots the true positive rate (the probability of detecting an anomaly) against the false positive rate (the probability of classifying a normal pixel as an anomaly) across all potential thresholds ($\tau$):
TPR $=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (26)
FPR $=\frac{\text{False Positives}}{\text{False Positives}+\text{True Negatives}}$ (27)
An algorithm’s performance improves as its ROC curve nears the top left of the ROC plot, indicating a good trade-off between detection and false alarms. Average ROC curves better evaluate ERX’s performance over repeated experiments due to its inherent variation from sparse random projection. The variance of the model is captured by the standard deviation of the ROC curves.
The area under each ROC curve (AUC) summarises the performance in a single metric. The AUC ranges from 0 to 1, where a score of 1 indicates a perfect model, and 0.5 indicates a model that is randomly guessing. A score below 0.5 indicates that the model consistently misclassifies anomalies as background, operating contrary to the intended task. Two additional metrics are calculated for model comparison: $\text{AUC}_{\text{TD}}$ to isolate target detectability and $\text{AUC}_{\text{BS}}$ to assess background suppressibility [80]. These metrics are slightly modified from their original form to have scores between 0 and 1:
$\text{AUC}_{\text{TD}}$ $=\frac{\text{AUC}+\text{AUC}_{\text{TPR},\tau}}{2}$ (28)
$\text{AUC}_{\text{BS}}$ $=\frac{\text{AUC}-\text{AUC}_{\text{FPR},\tau}+1}{2}$ (29)
where $\text{AUC}_{\text{TPR},\tau}$ is the area under the true positive rate and thresholds curve, and $\text{AUC}_{\text{FPR},\tau}$ is the area under the false positive rate and thresholds curve.
IV-D Experiments
The first experiment compares the speed and scalability of all algorithms, while the second evaluates the adaptability and robustness by analysing the detection performance. An ablation study on ERX examines the impact of varying momentum and using exponentially moving statistics, and assesses the effect of dimensionality reduction using sparse random projection.
All of the experiments are conducted on an NVIDIA Jetson Xavier NX device, a compact edge computer ideal for applications onboard small platforms such as a drone. This device features 8GB of RAM, a 6-core Carmel ARMv8.2 CPU, and is selected to operate in the 20W 6-core power mode. Real-time anomaly detection is simulated in a Python environment using a sampler and a detector. The sampler sequentially feeds each line (from left to right in the HSI datasets) to the detector (the algorithm being tested).
IV-D1 Speed Test